OpenStudy (anonymous):

[10.02]A 10−gallon container of water costs (y + 2) dollars while a 5−gallon container of water costs (y − 2) dollars. If the price per gallon of water is the same for both containers, the equation below can be used to solve for y. All of y plus 2 divided by 10 equals all of y minus 2 divided by 5 What is the value, in dollars, of y? (Do not put a $ in your answer.)

OpenStudy (anonymous):

right

OpenStudy (anonymous):

find the value of y on your own

OpenStudy (anonymous):

ummm ok

OpenStudy (anonymous):

is it 6?

OpenStudy (cruffo):

Translate: "All of y plus 2 divided by 10 equals all of y minus 2 divided by 5" \[\frac{y+2}{10} = \frac{y-2}{5}\] To solve equations containing fractions, multiply ever term on both sides of the equation by the LCD.

OpenStudy (anonymous):

is it 6?

OpenStudy (cruffo):

Is \[\frac{8}{10} = \frac{4}{5}?\]

OpenStudy (anonymous):

ummm ik sorry

OpenStudy (anonymous):

idk*

OpenStudy (cruffo):

\[\frac{y+2}{10} = \frac{y-2}{5}\] to check if 6 is a solution, replace the variable y with 6 to get \[\frac{8}{10} = \frac{4}{5}\] Note that \[\frac{8}{10} = \frac{8\div 2}{10 \div 2} = \frac{4}{5}\] So 6 is a good solution.